Assume that the average firm in your company's industry is expected to grow at a constant rate of 4% and that its dividend yield is 5%. Your company is about as risky as the average firm in the industry and just paid a dividend (D0) of $1.5. You expect that the growth rate of dividends will be 50% during the first year (g0,1 = 50%) and 20% during the second year (g1,2 = 20%). After Year 2, dividend growth will be constant at 4%. What is the estimated value per share of your firm’s stock? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Answer:

Required return = dividend yield + constant growth rate = 5% + 4% = 9%

D0 = 1.50

D1 = 1.50 * (1 + 50%) = 2.25

D2 = 2.25 * (1 +20%) = 2.70

D3 = 2.70 * (1 + 4%) = 2.808

P2 = D3 / (required return - growth rate) = 2.808 / (9% - 4%) = 56.16

Price = D1 / (1+return) + D2 / (1+return)^2 + P2 / (1+return)^2

Stock Price = 2.25 / 1.09 + 2.70 / 1.09^2 + 56.16 / 1.09^2 = 51.61

Answer: Stock price = $ 51.61